# Training-Conditional Coverage Bounds under Covariate Shift

### v1

@@ Appendix A Technical Lemmas: Weighted DKW Inequalities
$\operatorname{\mathbb{P}}\Bigg{(}\sup_{x\in\mathbb{R}}\left|\hat{F}_{Q,n}(x)-F_{Q}(x)\right|>\delta+3\,C\sqrt{\frac{B}{n}}\Bigg{)}\leq 4e^{-n\delta^{2}/(2B^{2})}.$
@@
$\int_{0}^{1}\sqrt{1+\log(2+1/\epsilon^{2})}\leq 1+\left(\int_{0}^{1}\log(2+1/\epsilon^{2})\right)^{1/2}\leq 1+\left(2\int_{0}^{1}(\epsilon^{2}-\log\epsilon)\right)^{1/2}\leq 3.$
Now let $P^{\prime}_{n}$ denote the empirical measure obtained after replacing $Z_{i}$ by some arbitrary $Z_{i}^{\prime}$. In this case, we observe that
$\Bigg{|}\sup_{x\in\mathbb{R}}\left|\int f_{x}~{}d(P_{n}-P)\right|-\sup_{x\in\mathbb{R}}\left|\int f_{x}~{}d(P_{n}^{\prime}-P)\right|\Bigg{|}$ $\leq\sup_{x\in\mathbb{R}}\left|\int f_{x}~{}d(P_{n}-P_{n}^{\prime})\right|$
@@
$\operatorname{\mathbb{E}}\left(\sup_{x\in\mathbb{R}}\left|\int f_{x}~{}d(P_{n}-P)\right|\right)\leq\frac{3\,CK}{\sqrt{n}}=\frac{C^{\prime}K}{\sqrt{n}}.$
@@
Lemma 3 (Alternative version)
@@ Appendix B Proof of Theorem 1
Recall that
$F_{Q,n}(x)=w_{\text{test}}\boldsymbol{1}\{t=\infty\}+\sum_{i\in\mathcal{I}}{w}_{i}\boldsymbol{1}\{s(Z_{i})\leq x\},\qquad{w}_{i}=\frac{\frac{dQ}{dP}(Z_{i})}{\frac{dQ}{dP}(Z_{\text{test}})+\sum_{i\in\mathcal{I}}\frac{dQ}{dP}(Z_{i})},$ (24)
@@
$P_{e}(\mathcal{D}_{n})=\operatorname{\mathbb{P}}\left(Y_{\text{test}}\not\in\hat{C}^{\text{split}}_{\alpha}(X_{\text{test}})\Big{|}\mathcal{D}_{n}\right)$ $=\operatorname{\mathbb{P}}\left(s(X_{\text{test}},Y_{\text{test}};\hat{\mu})>{F}^{-1}_{Q,m}(1-\alpha)\Big{|}\mathcal{D}_{n}\right)$
$\leq\operatorname{\mathbb{P}}\left(s(X_{\text{test}},Y_{\text{test}};\hat{\mu})>\hat{F}^{-1}_{Q,m}(1-\alpha)\Big{|}\mathcal{D}_{n}\right)$
@@
where $F_{Q}(x)=\operatorname{\mathbb{P}}_{Z\sim Q}(s(Z)\leq x|\mathcal{D}_{n})$ and $\hat{F}^{-1}_{Q,m}$ is defined according to (14) with index set $\mathcal{I}$ instead of $[n]$. Using the weighted DKW inequality from Lemma 1 we get
@@
$\operatorname{\mathbb{P}}\left(P_{e}(\mathcal{D}_{n})>\alpha+\delta+C^{\prime}\sqrt{\frac{B}{m}}\right)\leq 4e^{-m\delta^{2}/(2B^{2})}$
@@
$\operatorname{\mathbb{P}}\left(P_{e}(\mathcal{D}_{n})>\alpha+\sqrt{\frac{2B^{2}}{m}\log\frac{4}{\delta}}+C^{\prime}\sqrt{\frac{B}{m}}\right)$ $=\operatorname{\mathbb{P}}\left(P_{e}(\mathcal{D}_{n})>\alpha+\left(\sqrt{2B\log{4}/{\delta}}+C^{\prime}\right)\sqrt{\frac{B}{m}}\right)$
@@ B.1 Proof of Remark 1
$\operatorname{\mathbb{P}}\left(P_{e}(\mathcal{D}_{n})>\alpha+\delta\right)\leq\frac{6\,CK}{\delta\sqrt{m}}+\frac{4(K^{2}-1)}{m\delta^{2}}$
@@
$\operatorname{\mathbb{P}}\left(P_{e}(\mathcal{D}_{n})>\alpha+2\left(\frac{3CK+\sqrt{\delta(K^{2}-1)}}{\delta\sqrt{m}}\right)\right)\leq\delta$
@@ C.1 Proof of Theorem 2
$P_{e}({\bf D}_{n})=\operatorname{\mathbb{P}}(Y_{n+1}\notin\hat{\mathcal{C}}_{\alpha}^{\text{J+}}(X_{\text{test}})|{\bf D}_{n})$ $\leq\operatorname{\mathbb{P}}\left(1-F^{(n-1)}\left(\left|Y_{n+1}-\mu_{\overline{\beta}_{-1}}(X_{\text{test}})\right|\right)\leq\alpha+\delta+2\epsilon L_{n-1}\right)$

### v2

@@ Appendix A Technical Lemmas: Weighted DKW Inequalities
$\operatorname{\mathbb{P}}\Bigg{(}\sup_{x\in\mathbb{R}}\left|\hat{F}_{Q,n}(x)-F_{Q}(x)\right|>\delta+2C\sqrt{\frac{B}{n}}\Bigg{)}\leq 4e^{-n\delta^{2}/(2B^{2})}.$
@@
$\int_{0}^{1}\sqrt{1+\log(2+1/\epsilon^{2})}\leq\left(1+\int_{0}^{1}\log(2+1/\epsilon^{2})\right)^{1/2}\leq\left(1+2\int_{0}^{1}(\epsilon^{2}-\log\epsilon)\right)^{1/2}\leq 2.$
Now let $P^{(i)}_{n}$ denote the empirical measure obtained after replacing $Z_{i}$ by some arbitrary $Z_{i}^{\prime}$. In this case, we observe that
$\Bigg{|}\sup_{x\in\mathbb{R}}\left|\int f_{x}~{}d(P_{n}-P)\right|-\sup_{x\in\mathbb{R}}\left|\int f_{x}~{}d(P^{(i)}_{n}-P)\right|\Bigg{|}$ $\leq\sup_{x\in\mathbb{R}}\left|\int f_{x}~{}d(P_{n}-P^{(i)}_{n})\right|$
@@
$\operatorname{\mathbb{E}}\left(\sup_{x\in\mathbb{R}}\left|\int f_{x}~{}d(P_{n}-P)\right|\right)\leq\frac{2CK}{\sqrt{n}}=\frac{C^{\prime}K}{\sqrt{n}}.$
@@
Lemma 3 (Alternative version: bounded likelihood ratio)
@@
Lemma 4 (Alternative version: unbounded likelihood ratio)
Fix $B,\delta>0$ and $\epsilon\in(0,1/2)$. It holds that,
$\sup_{x\in\mathbb{R}}\left|F_{Q}(x)-\hat{F}_{Q,n}(x)\right|\leq 3\delta+4\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})>n\right)+C^{\prime}\sqrt{\frac{B}{n^{\prime}}}+\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})>B\right),$
with probability,
$1-4e^{-n^{\prime}\delta^{2}/(2B^{2})}$ $-2e^{-2n^{2\epsilon}}-2\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})>n\right)$
$-\frac{8}{n\delta^{2}}{\operatorname{\mathbb{E}}\left(\left(\frac{dQ}{dP}\right)^{2}(Z_{1})\boldsymbol{1}\left\{\frac{dQ}{dP}(Z_{1})\leq n\right\}\right)},$
where $n^{\prime}=n\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})\leq B\right)-n^{1/2+\epsilon}$.
Proof: To drop the assumption on $dQ/dP$, we start by writing $\mathcal{X}=\mathcal{X}_{B}\cup\mathcal{X}_{B}^{c}$, where $\mathcal{X}_{B}=\{x:dQ_{X}/dP_{X}(x)\leq B\}$. Define
$F_{Q|\mathcal{X}_{B}}:=\operatorname{\mathbb{P}}\left(s(Z_{\text{test}})\leq x\Big{|}\frac{dQ}{dP}(Z_{\text{test}})\leq B\right),$
$\hat{F}_{Q,n|\mathcal{X}_{B}}:=\frac{\sum_{i\in\mathcal{I}}\frac{dQ}{dP}(Z_{i})\boldsymbol{1}\{s(Z_{i})\leq x\}\boldsymbol{1}\left\{\frac{dQ}{dP}(Z_{i})\leq B\right\}}{\sum_{i\in\mathcal{I}}\frac{dQ}{dP}(Z_{i})\boldsymbol{1}\left\{\frac{dQ}{dP}(Z_{i})\leq B\right\}},$
where the summations are taken over calibration data points generated from $P$. We note,
$\sup_{x\in\mathbb{R}}\left|F_{Q}(x)-\hat{F}_{Q,n}(x)\right|\leq\sup_{x\in\mathbb{R}}\left|F_{Q}(x)-F_{Q|\mathcal{X}_{B}}(x)\right|$ $+\sup_{x\in\mathbb{R}}\left|F_{Q|\mathcal{X}_{B}}(x)-\hat{F}_{Q,n|\mathcal{X}_{B}}(x)\right|$
$+\sup_{x\in\mathbb{R}}\left|\hat{F}_{Q,n|\mathcal{X}_{B}}(x)-\hat{F}_{Q,n}(x)\right|.$
To upper-bound the first term, we observe that
$\left|F_{Q}(x)-F_{Q|\mathcal{X}_{B}}(x)\right|$ $=\left|\operatorname{\mathbb{P}}\left(s(Z_{\text{test}})\leq x\right)-\operatorname{\mathbb{P}}\left(s(Z_{\text{test}})\leq x\Big{|}\frac{dQ}{dP}(Z_{\text{test}})\leq B\right)\right|$
[…]
$+\operatorname{\mathbb{P}}\left(s(Z_{\text{test}})\leq x\Big{|}\frac{dQ}{dP}(Z_{\text{test}})>B\right)\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})>B\right)\Bigg{|}$
$\leq\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})>B\right)\Bigg{|}\operatorname{\mathbb{P}}\left(s(Z_{\text{test}})\leq x\Big{|}\frac{dQ}{dP}(Z_{\text{test}})>B\right)$
$\hskip 100.00015pt-\operatorname{\mathbb{P}}\left(s(Z_{\text{test}})\leq x\Big{|}\frac{dQ}{dP}(Z_{\text{test}})\leq B\right)\Bigg{|}$
$\leq\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})>B\right)$
Regarding the second term, using the same argument as in the proof of Lemma 1, we get
$\operatorname{\mathbb{P}}\bigg{(}\sup_{x\in\mathbb{R}}\left|F_{Q|\mathcal{X}_{B}}(x)-\hat{F}_{Q,n|\mathcal{X}_{B}}(x)\right|$ $\leq\delta+C^{\prime}\sqrt{B/n_{B}}\,\Big{|}\,n_{B}\bigg{)}\geq 1-4e^{-n_{B}\delta^{2}/(2B^{2})}$
where $n_{B}=\sum_{i\in\mathcal{I}}\boldsymbol{1}\left\{\frac{dQ}{dP}(Z_{i})\leq B\right\}$. Now let $n^{\prime}=n\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})\leq B\right)-n^{1/2+\epsilon}$ and observe that
$\operatorname{\mathbb{P}}$ $\bigg{(}\sup_{x\in\mathbb{R}}\left|F_{Q|\mathcal{X}_{B}}(x)-\hat{F}_{Q,n|\mathcal{X}_{B}}(x)\right|\leq\delta+C^{\prime}\sqrt{B/n^{\prime}}\bigg{)}$
$\geq\operatorname{\mathbb{P}}\bigg{(}\sup_{x\in\mathbb{R}}\left|F_{Q|\mathcal{X}_{B}}(x)-\hat{F}_{Q,n|\mathcal{X}_{B}}(x)\right|\leq\delta+C^{\prime}\sqrt{B/n_{B}},\,n_{B}\geq n^{\prime}\bigg{)}$
$=\operatorname{\mathbb{E}}\left(\operatorname{\mathbb{P}}\bigg{(}\sup_{x\in\mathbb{R}}\left|F_{Q|\mathcal{X}_{B}}(x)-\hat{F}_{Q,n|\mathcal{X}_{B}}(x)\right|\leq\delta+C^{\prime}\sqrt{B/n_{B}}\Big{|}n_{B}\bigg{)}\boldsymbol{1}\{n_{B}\geq n^{\prime}\}\right)$
$=\operatorname{\mathbb{E}}\left(\left(1-4e^{-n_{B}\delta^{2}/(2B^{2})}\right)\boldsymbol{1}\{n_{B}\geq n^{\prime}\}\right)$
$\geq\left(1-4e^{-n^{\prime}\delta^{2}/(2B^{2})}\right)\operatorname{\mathbb{P}}\left(n_{B}\geq n^{\prime}\right)\geq\left(1-4e^{-n^{\prime}\delta^{2}/(2B^{2})}\right)\left(1-2e^{-2n^{2\epsilon}}\right)$
$\geq 1-4e^{-n^{\prime}\delta^{2}/(2B^{2})}-2e^{-2n^{2\epsilon}}.$
For the third term, we note
$|{\hat{F}}_{Q,m|\mathcal{X}_{B}}(x)-{\hat{F}}_{Q,m}(x)|$ $=\left|\frac{W^{<}(x)}{W^{<}}-\frac{W^{<}(x)+W^{>}(x)}{W^{<}+W^{>}}\right|$
$=\left|\frac{W^{<}(x)W^{>}-W^{<}W^{>}(x)}{W^{<}(W^{<}+W^{>})}\right|$
$\leq 2\ \frac{W^{<}W^{>}}{W^{<}(W^{<}+W^{>})}=2\ \frac{W^{>}}{W^{<}+W^{>}}$ (24)
where
$W^{<}(x)=\sum_{i\in\mathcal{I}}\frac{dQ}{dP}(Z_{i})\boldsymbol{1}\{s(Z_{i})\leq x\}\boldsymbol{1}\left\{\frac{dQ}{dP}(Z_{i})\leq B\right\},$
$W^{>}(x)=\sum_{i\in\mathcal{I}}\frac{dQ}{dP}(Z_{i})\boldsymbol{1}\{s(Z_{i})\leq x\}\boldsymbol{1}\left\{\frac{dQ}{dP}(Z_{i})>B\right\},$
$W^{<}=\sum_{i\in\mathcal{I}}\frac{dQ}{dP}(Z_{i})\boldsymbol{1}\left\{\frac{dQ}{dP}(Z_{i})\leq B\right\},$
$W^{>}=\sum_{i\in\mathcal{I}}\frac{dQ}{dP}(Z_{i})\boldsymbol{1}\left\{\frac{dQ}{dP}(Z_{i})>B\right\},$
and the inequalty 24 follows from the fact that $W^{<}(x)\leq W^{<}$ and $W^{>}(x)\leq W^{>}$ for all $x\in\mathbb{R}$. Hence,
$\underset{x\in\mathbb{R}}{\sup}\ |{\hat{F}}_{Q,m|\mathcal{X}_{B}}(x)-{\hat{F}}_{Q,m}(x)|\leq 2\ \frac{\sum_{i\in\mathcal{I}}\frac{dQ}{dP}(Z_{i})\boldsymbol{1}\left\{\frac{dQ}{dP}(Z_{i})>B\right\}}{\sum_{i\in\mathcal{I}}\frac{dQ}{dP}(Z_{i})}\overset{a.s.}{\to}2\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})>B\right).$
We note,
$\left|\frac{\frac{1}{n}\sum_{i\in\mathcal{I}}\frac{dQ}{dP}(Z_{i})\boldsymbol{1}\left\{\frac{dQ}{dP}(Z_{i})>B\right\}}{\frac{1}{n}\sum_{i\in\mathcal{I}}\frac{dQ}{dP}(Z_{i})}-\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})>B\right)\right|\leq$
$\left|{\frac{1}{n}\sum_{i\in\mathcal{I}}\frac{dQ}{dP}(Z_{i})\boldsymbol{1}\left\{\frac{dQ}{dP}(Z_{i})>B\right\}}-\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})>B\right)\right|+$
$\hskip 50.00008pt\left|\frac{1}{{\frac{1}{n}\sum_{i\in\mathcal{I}}\frac{dQ}{dP}(Z_{i})}}-1\right|\left|{\frac{1}{n}\sum_{i\in\mathcal{I}}\frac{dQ}{dP}(Z_{i})\boldsymbol{1}\left\{\frac{dQ}{dP}(Z_{i})>B\right\}}\right|\leq$
$\left|{\frac{1}{n}\sum_{i\in\mathcal{I}}\frac{dQ}{dP}(Z_{i})\boldsymbol{1}\left\{\frac{dQ}{dP}(Z_{i})>B\right\}}-\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})>B\right)\right|+\left|{\frac{1}{n}\sum_{i\in\mathcal{I}}\frac{dQ}{dP}(Z_{i})}-1\right|\leq$
$\left|{\frac{1}{n}\sum_{i\in\mathcal{I}}\frac{dQ}{dP}(Z_{i})\boldsymbol{1}\left\{\frac{dQ}{dP}(Z_{i})>B\right\}}-\operatorname{\mathbb{P}}\left(B<\frac{dQ}{dP}(Z_{\text{test}})\leq n\right)\right|+$
$\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})>B,\frac{dQ}{dP}(Z_{\text{test}})>n\right)+\left|{\frac{1}{n}\sum_{i\in\mathcal{I}}\frac{dQ}{dP}(Z_{i})}-\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})\leq n\right)\right|+\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})>n\right)$
$\leq\delta+2\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})>n\right)$
with probability at least Durrett [2019][Theorem 2.2.12],
$1-2\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})>n\right)-\frac{8}{n\delta^{2}}{\operatorname{\mathbb{E}}\left(\left(\frac{dQ}{dP}\right)^{2}(Z_{1})\boldsymbol{1}\left\{\frac{dQ}{dP}(Z_{1})\leq n\right\}\right)}.$
Thus,
$\sup_{x\in\mathbb{R}}\left|F_{Q}(x)-\hat{F}_{Q,n}(x)\right|\leq 3\delta+4\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})>n\right)+C^{\prime}\sqrt{\frac{B}{n^{\prime}}}+\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})>B\right),$
with probability at least,
$1-4e^{-n^{\prime}\delta^{2}/(2B^{2})}-2e^{-2n^{2\epsilon}}-2\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})>n\right)-\frac{8}{n\delta^{2}}{\operatorname{\mathbb{E}}\left(\left(\frac{dQ}{dP}\right)^{2}(Z_{1})\boldsymbol{1}\left\{\frac{dQ}{dP}(Z_{1})\leq n\right\}\right)}.$
Remark 2
Observe that ${\hat{F}}_{Q,m|\mathcal{X}_{B}}(x)-{\hat{F}}_{Q,m}(x)\overset{a.s.}{\to}{F}_{Q|\mathcal{X}_{B}}(x)-{F}_{Q}(x)$. Also note that ${\hat{F}}_{Q,m|\mathcal{X}_{B}}(x)-{\hat{F}}_{Q,m}(x)\overset{a.s.}{\to}0$ if $s(Z_{\text{test}})$ and $\frac{dQ}{dP}(Z_{\text{test}})$ are independent.
@@ Appendix B Proof of Theorem 1
Recall from Section 2.3
$F_{Q,|\mathcal{I}|}(x)=w_{\text{test}}\boldsymbol{1}\{t=\infty\}+\sum_{i\in\mathcal{I}}{w}_{i}\boldsymbol{1}\{s(Z_{i})\leq x\},\qquad{w}_{i}=\frac{\frac{dQ}{dP}(Z_{i})}{\frac{dQ}{dP}(Z_{\text{test}})+\sum_{i\in\mathcal{I}}\frac{dQ}{dP}(Z_{i})},$ (25)
@@
$P_{e}({\bf D}_{n})=\operatorname{\mathbb{P}}\left(Y_{\text{test}}\not\in\hat{C}^{\text{split}}_{\alpha}(X_{\text{test}})\Big{|}{\bf D}_{n}\right)$ $=\operatorname{\mathbb{P}}\left(s(X_{\text{test}},Y_{\text{test}};\hat{\mu})>{F}^{-1}_{Q,m}(1-\alpha)\Big{|}{\bf D}_{n}\right)$
$\leq\operatorname{\mathbb{P}}\left(s(X_{\text{test}},Y_{\text{test}};\hat{\mu})>\hat{F}^{-1}_{Q,m}(1-\alpha)\Big{|}{\bf D}_{n}\right)$
@@
where $F_{Q}(x)=\operatorname{\mathbb{P}}_{Z\sim Q}(s(Z)\leq x)$ and $\hat{F}^{-1}_{Q,m}$ is defined according to (14) with index set $\mathcal{I}$ instead of $[n]$. Using the weighted DKW inequality from Lemma 1 we get
@@
$\operatorname{\mathbb{P}}\left(P_{e}({\bf D}_{n})>\alpha+\delta+C^{\prime}\sqrt{\frac{B}{m}}\right)\leq 4e^{-m\delta^{2}/(2B^{2})}$
@@
$\operatorname{\mathbb{P}}\left(P_{e}({\bf D}_{n})>\alpha+\sqrt{\frac{2B^{2}}{m}\log\frac{4}{\delta}}+C^{\prime}\sqrt{\frac{B}{m}}\right)$ $=\operatorname{\mathbb{P}}\left(P_{e}({\bf D}_{n})>\alpha+\left(\sqrt{2B\log{4}/{\delta}}+C^{\prime}\right)\sqrt{\frac{B}{m}}\right)$
@@ B.1 Proof of Remark 1
$\operatorname{\mathbb{P}}\left(P_{e}({\bf D}_{n})>\alpha+\delta\right)\leq\frac{6\,CK}{\delta\sqrt{m}}+\frac{4(K^{2}-1)}{m\delta^{2}}$
@@
$\operatorname{\mathbb{P}}\left(P_{e}({\bf D}_{n})>\alpha+2\left(\frac{2CK+\sqrt{\delta(K^{2}-1)}}{\delta\sqrt{m}}\right)\right)\leq\delta$
@@ C.1 Proof of Theorem 3
$P_{e}({\bf D}_{n})=\operatorname{\mathbb{P}}(Y_{\text{test}}\notin\hat{\mathcal{C}}_{\alpha}^{\text{J+}}(X_{\text{test}})|{\bf D}_{n})$ $\leq\operatorname{\mathbb{P}}\left(1-F^{(n-1)}\left(\left|Y_{\text{test}}-\mu_{\overline{\beta}_{-1}}(X_{\text{test}})\right|\right)\leq\alpha+\delta+2\epsilon L_{n-1}\right)$